### v2

Abstract
We have discovered the stellar counterpart to the ALFALFA Virgo 7 cloud complex, which has been thought to be optically dark and nearly star-free since its discovery in 2007. This 190 kpc long chain of enormous atomic gas clouds ( M) is embedded in the hot intracluster medium of the Virgo galaxy cluster but is isolated from any galaxy. Its faint, blue stellar counterpart, BC6, was identified in a visual search of archival optical and UV imaging. Follow-up observations with the Green Bank Telescope, Hobby-Eberly Telescope, and Hubble Space Telescope demonstrate that this faint counterpart is at the same velocity as the atomic gas, is actively forming stars, and is metal-rich (). We estimate its stellar mass to be only , making it one of the most gas-rich stellar systems known. Aside from its extraordinary gas content, the properties of BC6 are entirely consistent with those of a recently identified class of young, low-mass, isolated, and star-forming clouds in Virgo, that appear to have formed via extreme ram pressure stripping events. We expand the existing discussion of the origin of this structure and suggest NGC 4522 as a likely candidate, however, the current evidence is not fully consistent with any of our proposed progenitor galaxies. We anticipate that other “dark” gas clouds in Virgo may have similarly faint, star-forming counterparts. We aim to identify these through the help of an ongoing citizen science search of the entire cluster.
1 Introduction
Surveys of atomic gas (H i) have a long history of identifying apparently dark structures made up of large quantities of neutral gas and (almost) no stars (e.g. Giovanelli & Haynes, 1989; Chengalur et al., 1995; Verdes-Montenegro et al., 2001; Davies et al., 2004; Minchin et al., 2005; Haynes et al., 2007; Kent et al., 2007; Taylor et al., 2012; Wong et al., 2021; Józsa et al., 2022; Jones et al., 2023). H i is prone to forming such structures as it is typically the most loosely bound baryonic component of a galaxy’s disk and is thus the most easily removed by external forces. Furthermore, as the distribution of H i often extends well beyond the stellar disk of a galaxy it is relatively straightforward to strip gas, but no stars, thereby producing optically dark clouds.
Although some of the structures identified in the references above are still connected to their parent galaxies, some are detached, isolated, and apparently dark H i clouds with no clear point of origin. Such objects span several orders of magnitude in H i mass and are certainly not a single type of object. At the high-mass end, Kent et al. (2007) and Taylor et al. (2012) identified several dark H i clouds () in the direction of the Virgo cluster. At intermediate masses, the Five-hundred-meter Aperture Spherical Telescope and the Green Bank Telescope (GBT) have recently identified lone H i clouds that may be ancient objects (Zhou et al., 2023; Karunakaran & Spekkens, 2024, K. O’Neil private communication). At the low-mass end, Adams et al. (2013) identified a sample of optically dark, compact H i clouds, some of which could be very low mass halos in or near the Local Group.
When dealing with H i-only detections of local objects (e.g. km s) it is challenging to determine accurate distances, and these systems may not be as they first appear. Adams et al. (2015), Bellazzini et al. (2015), and Sand et al. (2015) all, almost simultaneously, realized that one of the Adams et al. (2013) objects (AGC 226067) did contain a small number of stars and was 10 farther away and 100 more massive than first thought. This object, now commonly referred to as SECCO 1, was found to be an isolated pocket of recent star formation (SF) embedded in the hot intra-cluster medium (ICM) of the Virgo cluster. Its rich metallicity (Beccari et al., 2017) and exclusively young stellar population (Sand et al., 2017) indicated that it must have formed recently from stripped gas, but there was no candidate parent galaxy within over 200 kpc.
Sand et al. (2015) searched for other objects with similar optical and UV morphology to SECCO 1, identifying five additional candidates in Virgo. With Hubble Space Telescope (HST) imaging, plus Jansky Very Large Array (VLA), GBT, and Multi Unit Spectroscopic Explorer (MUSE) observations (Jones et al., 2022a, b; Bellazzini et al., 2022) we demonstrated that four222The fifth object is resolved as a background galaxy group in HST imaging (Jones et al., 2022b). of these are analogous to SECCO 1 (though not all are H i-rich) and argued that they represent a new class of object, which we will refer to simply as “blue blobs.” The only viable mechanism to explain both the isolation and young ages of blue blobs is high speed ram pressure stripping of H i-bearing galaxies falling into the cluster (Jones et al., 2022b). Even so, it is remarkable that they survived traversing 100s of kpc through the hot ICM, and may be the first examples of isolated star-forming clouds supported (in part) by external pressure (Burkhart & Loeb, 2016; Bellazzini et al., 2018; Calura et al., 2020; Jones et al., 2022b).
With the goal of better understanding the population statistics of these new objects we embarked on a uniform search for candidates covering the entire cluster (described in §2). When we began following up our initial candidates (§3), one (known as BC6) was found to be coincident (both in position and velocity) with the largest of the “dark” H i cloud complexes identified by Kent et al. (2007). Despite repeated study of this cloud complex (Kent et al., 2009; Sorgho et al., 2017; Minchin et al., 2019) no convincing stellar counterpart had been identified since its discovery. Here we show (§4) that it has a definite stellar counterpart that is over 20,000 times less massive than the H i, and is made up exclusively of stars that recently formed in situ. The stellar properties of this object closely match those of the recently identified blue blobs (Jones et al., 2022b). BC6 differs only in its exceptional H i content. In §5 we discuss the potential point of origin of this object, and its likely fate. We present our conclusions in §6.
Throughout this letter we attempt to use “BC6” to refer to the stellar component which we have identified, and the “H i cloud complex” to refer to the (previously dark) H i structure identified by Kent et al. (2007). We also assume throughout that BC6 and the associated H i complex are in the Virgo cluster at a distance of 16.5 Mpc (Mei et al., 2007). We note that this assumption is consistent with all observations, in particular the color–magnitude diagram (§3.4 & §4.1).
2 Identification of candidates
The blue blob candidates presented in Jones et al. (2022b) arose from a cursory search (Sand et al., 2015) of the Virgo cluster and there are likely additional examples. In an attempt to identify additional objects we began a visual search of the available imaging within the cluster from the Next Generation Virgo cluster Survey (NGVS, Ferrarese et al., 2012), the Dark Energy Camera Legacy Survey (DECaLS Dey et al., 2019), and the Galaxy Evolution Explorer (GALEX Martin et al., 2005). A full description of this search will be presented in Dey et al. (in prep.); here we provide a brief summary.
Matching 3′3′ image cutouts from the three surveys were produced, covering the full area of the NGVS. The DECaLS color images were taken directly from the Legacy Surveys Viewer333www.legacysurvey.org/viewer, as were the GALEX NUV+FUV images. In the case of NGVS, we produced our own RGB images from the , and filters, with a stretch designed to bring out faint, blue features (saturating most normal galaxies in the cluster). These image cutouts were uploaded to the Zooniverse website where team members (primarily M. Jones, S. Dey, and D. Sand) searched for objects that had UV emission (or were exceptionally blue in the optical), but had extremely faint, irregular, and clumpy appearances in the optical images.444This initial search has since evolved into a full citizen science project that will be presented in detail in Mazziotti et al. (in prep) and Dey et al. (in prep). It is available at https://www.zooniverse.org/projects/mike-dot-jones-dot-astro/blobs-and-blurs-extreme-galaxies-in-clusters.
The most compelling blue blob candidates to emerge from this search, one of which was Blue Candidate 6 (BC6, Figure 1), were selected for follow-up observations with GBT, HST, and HET, to search for H i gas content, resolve their stellar populations, and search for optical emission lines, respectively. These observations are discussed in §3.
3 Observations
BC6 lies in the direction of the Virgo cluster approximately 2.7 south of M 87 and 1.7 north of M 49. That is, roughly halfway been the Virgo A and B clouds (Binggeli et al., 1987). Typical cluster members in this direction have line of sight velocities in the range , but the scatter is several hundred km s beyond either side of this range (Mei et al., 2007). There is a bridge of diffuse X-ray emission that connects M 87 to M 49 (Böhringer et al., 1994), indicating that the hot ICM is higher density in this region than would be expected given its separation from the cluster core. BC6 lies roughly on this bridge.
3.1 NGVS, DECaLS, and GALEX
BC6 appears to have three main components (Figure 1). The largest and brightest is just south of a background galaxy that is visible in both optical and UV images. This main component (BC6a) contains a bright knot of FUV emission (blue in Figure 1, right) and appears as a faint, clumpy, blue smudge in the NGVS image. The other two components are even fainter and are to the NE and east of the brightest component. BC6b (to the NE) has faint, but still clearly discernible UV emission, while BC6c (to the east) is barely detected in UV and appears only as a few blue points in the NGVS image.
Aperture photometry was performed on the NGVS and filters and GALEX NUV using the Aperture Photometry Tool (Laher et al., 2012). Due to the irregular morphology of BC6, elliptical apertures were manually tailored for each component and the median sky level subtracted based on annuli beyond them. Galactic extinction was corrected using the dust maps of Schlegel et al. (1998) and the filter coefficients from Schlafly & Finkbeiner (2011) and Wyder et al. (2007). Using standard conversions (Iglesias-Páramo et al., 2006; Morrissey et al., 2007) we converted the GALEX NUV flux into a star formation rate (SFR) estimate. These photometry values are listed in Table 1.
3.2 Green Bank Telescope
After the identification of BC6 and several other candidates in the NGVS and GALEX imaging, we rapidly embarked on a follow-up program with a joint GBT-HST project, GBT-22B-080 (PI: M. Jones), to search for H i gas in all candidates and to obtain high resolution imaging of the stellar population of BC6. In addition, we followed up candidates with a Hobby-Eberly Telescope (HET) project (PI: S. Janowiecki) to obtain optical spectroscopy of nebular emission lines (§3.5).
Before it was realized that BC6 was the stellar counterpart to the H i structure identified by Kent et al. (2007), and that it therefore already had H i observations, it was included as a target in the GBT project. BC6 was observed in April 2023 with six 10 min on/off integrations with the VEGAS (Versatile GBT Astronomical Spectrometer) backend. The data were reduced using standard procedures in GBTIDL.555https://gbtidl.nrao.edu After smoothing to a spectral resolution of 5 km s, the resulting spectrum (Figure 2) has an rms noise of 1.77 mJy. H i line emission is strongly detected in the 9.1′ GBT beam along the BC6 line-of-sight between 475 km s – 625 km s.
3.3 ALFALFA
The Arecibo Legacy Fast ALFA (Arecibo L-band Feed Array), or ALFALFA, survey (Giovanelli et al., 2005; Haynes et al., 2011, 2018) mapped approximately 3000 sq deg of the sky for H i line emission, including the Virgo cluster. Kent et al. (2007) originally identified the H i cloud complex in a preliminary ALFALFA data set. These data have a channel width of 5 km s, a pixel size of 1′, and a spatial resolution of approximately 3.8′. We created our own version of the moment zero map of the H i emission in the cloud complex using the SoFiA (Serra et al., 2014, 2015) masking tool, adopting the standard smooth and clip algorithm with spatial smoothing kernels of 4.0 and 8.0′, with both no spectral smoothing and smoothing over two and four channels (10 and 20 km s). We set a detection threshold of 3.5 in order to include relatively faint emission, but coupled this with a high reliability threshold (95%) in order to ensure that only sources that are high S/N overall were retained. The moment zero map of the emission included in the SoFiA mask is shown in Figure 3 (left) as contours on a DECaLS optical image.
We estimate the H i mass of the entire structure by integrating all of the flux within the SoFiA mask and assuming a distance of 16.5 Mpc. This gives a value of . We note that this is about 0.1 dex higher than the value in Kent et al. (2009), however, in that work they considered each cloud component as a separate source, which may have led to some extended flux being missed. In addition, Minchin et al. (2019) used deeper Arecibo observations to estimate a total H i mass of , arguing that the additional flux (compared to ALFALFA) comes from low column density, extended emission.
3.4 Hubble Space Telescope imaging
HST Advanced Camera for Surveys (ACS) images of BC6666http://dx.doi.org/10.17909/h12e-gb42 (catalog DOI: 10.17909/h12e-gb42) were taken in February 2023 as part of the program GO-17267 (PI: M. Jones). The imaging used the Wide Field Channel of ACS with the F606W and F814W filters. Over the course of two orbits, four exposures were taken in each filter, resulting in total integration times of approximately 2000 s in F606W and F814W. A false color composite of these images is shown in Figure 3 (right), centered on BC6.
Point source photometry was carried out with DOLPHOT (Dolphin, 2000, 2016), which uses the Vega magnitude system. Thus, our point source photometry is in the Vega system. Quality cuts to eliminate unwanted sources (e.g. background galaxies and diffraction spikes) were made as described in Section 3.1 of Jones et al. (2022b). Artificial star tests were used to estimate the 90% and 50% completeness limits of the source catalog (as in Jones et al., 2022a). The Python package dustmaps (Green, 2018) was used to correct the point source photometry for Galactic extinction using the maps of Schlegel et al. (1998) and the coefficients from Schlafly & Finkbeiner (2011) for the F606W and F814W filters. The typical extinction was very minor with a mean E(B-V) value of 0.024 mag. Stars within the ellipses in Figure 3 (right) are shown in the CMD in Figure 4. A background CMD is also shown for an equal area region far from BC6.
In addition to the point source photometry we also performed aperture photometry on the HST images in order to measure the integrated magnitude of BC6 in the F606W and F814W filters. To do this we used the drizzled images in each filter and the apertures plotted in Figure 3 (right). Owing to the highly irregular morphology of BC6 these apertures were manually constructed to encompass the regions containing blue stars and UV emission, as well as with the distribution of light in the NGVS image (Figure 1, left). Using the Photutils (Bradley et al., 2020) package the integrated magnitude within these apertures was measured and the background contribution subtracted based on the median sky value within annular apertures around each component. For comparison purposes we also convert to absolute V-band magnitude based on Sirianni et al. (2005) and an assumed distance of 16.5 Mpc. These magnitude measurements are quoted in Table 1.
3.5 Hobby-Eberly Telescope optical spectroscopy
We observed the main component of BC6 (the largest ellipse in Figure 3, right) with the Blue spectrograph of the Low Resolution Spectrograph 2 (LRS2-B, Chonis et al., 2016) on the 10m HET (Ramsey et al., 1998; Hill et al., 2021) under clear dark skies on the night of April 19th, 2023 with a 3300 s exposure. The LRS2-B spectrograph has and an integral field unit which is 6″12″ with 0.6″ lenslet fibers on sky, giving an abundance of sky fibers around the relatively compact optical emission from BC6. We use the H line to determine which spaxels include emission from BC6 and extract all spaxels within 1.5″ of the peak of the source with the LRS2Multi extension to the Panacea data reduction pipeline.777https://github.com/grzeimann/LRS2Multi and https://github.com/grzeimann/Panacea Briefly, this data reduction includes the necessary detector characterizations, fiber extraction, wavelength calibration, mirror illumination correction, sky subtraction, and 1-D spectral source extraction including a treatment of differential atmospheric refraction in the sky and object spectra. The resulting spectrum is shown in Figure 5.
The spectrum is fairly typical of an H ii region, with clear, albeit relatively weak, emission lines. Only the H line is high enough S/N to be fit with a Gaussian. Thus, this line alone was used to determine the redshift km s (with a barycentric correction applied). To obtain fluxes from the other lines we extract the integrated signal within a 2 region centered on the redshifted wavelength on each line. Here is the standard deviation of the Gaussian profile fit to the H line. Flux errors are estimated by summing the uncertainty values in the same wavelength range and dividing by the square root of the number of channels included.
A reddening correction was calculated by assuming an intrinsic value of (for an H ii region at K; Osterbrock, 1989), an value of 3.1, and the Fitzpatrick (2004) extinction law. Based on the observed line ratio we calculated an extinction value of . This suggests that there is some internal extinction within BC6 as the Schlegel et al. (1998) dust map in this direction indicates minimal Galactic extinction (). Observed line fluxes were corrected for extinction (using the extinction model of Fitzpatrick, 2004) implemented in the Python package dust_extinction.888https://dust-extinction.readthedocs.io These fluxes and uncertainties are provided in Table 2. For consistency the quoted H flux does not correspond to that calculated with the Gaussian fit, but was instead calculated in the same manner as the other lines.
With these line flux measurements we used two line ratio metrics to estimate the gas-phase oxygen abundance in BC6 (following Pettini & Pagel, 2004), N2=[Nii]/H and O3N2=([Oiii]/H)/([Nii]/H). We note that these ratios are designed to be almost independent of our extinction correction as [Oiii] and H occur at almost the same wavelength, as do [Nii] and H. We see (in Table 2) that both metrics give almost exactly the same oxygen abundance, despite each method having an intrinsic scatter of 0.2 dex.
Finally, we made a second estimate of the SFR based on the H line flux. Following the standard Kennicutt (1998) conversion this equates to a SFR of .
4 Results
Parameter | Value |
RA (J2000) | 12:29:35 |
Dec. (J2000) | 9:44:32 |
/km s | |
/mag (AB) | |
/mag (AB) | |
/mag (AB) | |
/mag (AB) | |
/mag (AB) | |
6.05 | |
/M | 4.4 |
H | H | [Nii] | [Oiii] | ||
4.1 Morphology and stellar population
The optical appearance of BC6 is remarkably similar to that of BC4, which Jones et al. (2022b) concluded was likely in the process of becoming gravitationally unbound after losing its H i gas reservoir (presumably the majority of its initial total mass). Like BC4, BC6 is broken up into a few main components that are separated by 10s of arcsec (10″ pc at 16.5 Mpc). Each component contains multiple clumps of blue stars and has associated UV emission. In the HST image these clumps of blue stars are partially resolved, with the brightest stars individually detected.
In the CMD of these resolved stars (Figure 4) we see a very similar stellar population to those of other blue blobs (Sand et al., 2017; Jones et al., 2022a, b): a population of predominantly very blue (F606W-F814W = 0) stars, and a second population at approximately F606W-F814W = 1. As in our previous works, we argue that this population is consistent with very young main sequence stars and the blue and red sides of the Helium burning branch at the distance of the Virgo cluster (16.5 Mpc). The PARSEC (PAdova and TRieste Stellar Evolution Code; Bressan et al., 2012) isochrones overplotted in the middle panel of Figure 4 indicate that the entirety of the observed population is consistent with stars in the age range of 10-100 Myr. In particular, we point out that these isochrones were not fitted to the CMD in any way, yet the color range of the Helium burning stars in the isochrones matches closely with the observed color range. This range is mainly determined by the metallicity of the stars, which we independently measured (and input to the isochrones) from optical spectroscopy of the H ii regions (§3.5 & §4.2).
We also see no evidence of an old population in the form of a red giant branch (RGB). However, at the distance of Virgo the tip of the RGB should occur at approximately F814W = 27 (Jang & Lee, 2017), which is roughly the location of our 50% completeness limit. Even so, if a significant underlying old population were present, we would expect to see a build up of RGB stars around the completeness limit. The absence of this population implies that BC6’s stellar population is purely a young population, however, without deeper imaging it is impossible to determine the exact age of the oldest stars.
The two SFR estimates (Table 1) from NUV and H differ by almost an order of magnitude. Although the HET observation only covered a single component of BC6, it was the main body (Figure 3, right), where most of the UV flux originates. As NUV is sensitive to SF over a longer timescale than H, this may be an indication that SF is currently ramping down in BC6. However, given the very low SFR in BC6 it is also possible that this is merely indicative of stochasticity in the number of high mass stars produced in any given 10 Myr period.
4.2 Metallicity
Beccari et al. (2017) and Bellazzini et al. (2022) measured the metallicities of five blue blobs (SECCO1, BC1, and BC3-5) using the same line ratio metrics as used here (see §3.5) and found a range of . BC6 fits comfortably within this range (). Thus, as with its stellar population, it seems quite consistent with the established sample of blue blobs.
Following the same rationale as Jones et al. (2022b), the above metallicity estimate implies that the progenitor galaxy of BC6 was likely in the stellar mass range , assuming it roughly follows the mass–metallicity relation (Andrews & Martini, 2013). The broad range is a reflection of the large uncertainty in the metallicity. Unfortunately, over this stellar mass range the typical gas fraction () of a field galaxy smoothly transitions over the range (e.g. Huang et al., 2012), thus the fact that BC6 is accompanied by 10 M of H i gas does little to help narrow down the progenitor’s stellar mass. However, it does mean that the progenitor must have lost a significant fraction, probably the majority, of its H i gas reservoir.
4.3 Stellar mass
To estimate the stellar mass of BC6 we follow the same approach as Jones et al. (2022b), that is, assume a constant SFR and continuously build up a synthetic stellar population (of matching metallicity) until the observed total F814W magnitude of BC6 is reached. Using this approach we estimate a stellar mass of M, which at the observed (NUV-based) SFR could have been built up over approximately 80 Myr. If we instead use and -band aperture photometry from the NGVS images and the scaling relations of Zibetti et al. (2009) and Taylor et al. (2011), we obtain estimates of and M, respectively.
It is encouraging that all these estimates fall within a single order of magnitude, even if they still vary by a factor of five. Although our initial estimate is smaller than both the others, we expect that this is the most representative of the actual stellar mass of BC6. Fundamentally, all three estimates rely on a mass-to-light scaling to produce a stellar mass, but only in our initial estimate is the synthetic population used to generate that mass-to-light ratio based on the observed metallicity and current SFR of BC6. This population is also quite atypical, with no apparent old population (older than a few hundred Myr), thus standard scaling relations are unlikely to be entirely appropriate.
With a stellar mass of M, BC6 must have a gas fraction () of around 3,000 if we only consider the gas in the coincident cloud (C3; Figure 3, left), or over 20,000 if we include the whole H i cloud complex. This makes it one of, if not the most H i-rich extragalactic stellar systems known.
5 Discussion
5.1 Point of origin
Although the stellar counterpart, BC6, was unknown until now, the origin of the H i cloud complex has been the subject of considerable discussion in the literature since it was first detected (Kent et al., 2007, 2009; Sorgho et al., 2017; Minchin et al., 2019). However, the gas-phase metallicity offers a new piece of pertinent information, and we also approach the question of the point of origin from a new standpoint in light of similar discussions for other blue blobs (Adams et al., 2015; Beccari et al., 2017; Sand et al., 2017; Bellazzini et al., 2018; Calura et al., 2020; Jones et al., 2022a, b).
As mentioned by Minchin et al. (2019), tidal stripping does not seem to be a viable mechanism to form this structure. Not only is its isolation more difficult to explain with the low velocities (300 km s) involved in tidal stripping (e.g. Bournaud & Duc, 2006), but M represents a significant amount of H i gas for almost any galaxy and it is not possible to tidally remove a large fraction of a galaxy’s gas disk without also stripping stars. The exception to this would be if the progenitor was a particularly massive and H i-rich galaxy, however, that would also require a close passage from a similarly massive perturber to strip the gas, with neither remaining connected to the stripped gas in the present. Thus, the complete lack of stars over almost all of such a massive H i complex is a strong argument for ram pressure stripping and we will only consider the formation of BC6 via this mechanism.
The H i cloud complex (Figure 3) is about 40′ across (190 kpc), making it plausible that the progenitor galaxy is a degree or more away. Unfortunately, despite there being no clear progenitor galaxy in the immediate vicinity of BC6, there are 10s of potential parent objects that are Virgo members within a few degrees. Both Kent et al. (2009) and Minchin et al. (2019) focused on NGC 4445 and NGC 4424 as these are only 30′ to the SW of BC6 and have similar radial velocities to the gas complex (346 km s and 437 km s, respectively; Albareti et al., 2017; Haynes et al., 2018). NGC 4424 is in the process of being ram pressure stripped (Chung et al., 2009; Sorgho et al., 2017), but its H i tail points to the SE, roughly parallel to the H i complex containing BC6, rather than towards it. In the case of NGC 4445 the –B-band diameter scaling relation of Jones et al. (2018) indicates that its expected original H i content (before falling into the Virgo cluster) would have been (and it still contains ; Haynes et al., 2018). The observed mass of the H i complex therefore represents almost its entire original H i content. Furthermore, in Sorgho et al. (2017), figure A1, NGC 4445 also appears to have a ram pressure tail (though much smaller than that of NGC 4424) which is pointing almost directly away from BC6 and the H i cloud complex. Thus, neither of these appear to be a viable candidate for the progenitor system.
In simulations of high speed ram pressure stripping events (e.g. Kapferer et al., 2009; Tonnesen & Bryan, 2012, 2021; Göller et al., 2023), on scales much larger than a galaxy, the tails that are formed are usually relatively linear features. This means that for progenitor candidates that are within 1, but not aligned with the major axis of the H i complex, a larger than expected opening angle would be required to explain the geometry of the structure. It would perhaps still be possible to reproduce the observed geometry by involving clumpy structure within the ICM (the simulations assume the ICM is uniform and moving at a constant velocity). However, in an attempt to narrow our focus to the most likely candidates, we will now consider potential progenitors that are roughly NW and SE of BC6, such that the cloud geometry can readily be explained by a simple ram pressure stripping scenario. There are five potentially viable progenitors within a few degrees: NGCs 4380 and 4390 to the NW, and NGCs 4519, 4522, and 4535 to the SE. All five are massive enough to meet the mass–metallicity requirement from §4.2 and all currently contain some H i, detected in ALFALFA. The basic properties of these galaxies are summarized in Table 3.
Galaxy | /km s | /km s | Separation | Direction | [M/H] | ||
NGC 4380 | 963 | 463 | 64′ | NW | 8.51 | 9.40 | |
NGC 4390 | 1101 | 601 | 70′ | NW | 8.68 | 8.85 | -0.38 |
NGC 4424 | 437 | -63 | 40′ | W | 8.34 | 9.43 | 0.04 |
NGC 4445 | 353 | -147 | 27′ | W | 7.71 | 9.19 | |
NGC 4519 | 1216 | 716 | 87′ | SE | 9.47 | 9.33 | -0.18 |
NGC 4522 | 2329 | 1819 | 69′ | SE | 8.66 | 9.45 | -0.01 |
NGC 4535 | 1963 | 1463 | 116′ | SE | 9.57 | 9.93 | 0.06 |
We see from the metallicity measurements that NGC 4519 and NGC 4522 are the closest matches to BC6 (which has ). However, there is no measurement for NGC 4380, and NGC 4535 could still be a match given the uncertainties. Thus, only NGC 4390 is strongly disfavored because of its metallicity. The predicted original H i mass of NGC 4390 also suggests that it never would have had enough H i to form the cloud complex associated with BC6. Furthermore, it appears NGC 4390 has not lost a significant amount of H i.999We note that the predicted original H i mass is usually intended as a statistical quantity and has considerable uncertainty for any individual galaxy. For example, the measured H i content of NGC 4519 exceeds its predicted original H i mass by over 0.1 dex. These values should be considered a rough guide only.
NGC 4380 was mapped in the VLA Imaging of Virgo in Atomic gas project (VIVA; Chung et al., 2009), which found that it has a truncated H i disk. This indicates a past episode of ram pressure stripping, however, the direction of stripping is unknown.
Unfortunately, VIVA did not map either NGC 4519 or NGC 4535. Both are detected in ALFALFA, but these data have an angular resolution of 4′. At this resolution NGC 4535 appears undisturbed, but the H i distribution of NGC 4519 is clearly extended to the NW (towards BC6). However, NGC 4519 has a neighbor immediately to the NW (NGC 4519A) separated by only 200 km s (Binggeli et al., 1985). Thus the disturbance in the H i distribution of NGC 4519 may be the result of either tides or ram pressure and with the resolution of the ALFALFA data there is little chance of distinguishing the two.
NGC 4522 was mapped in VIVA and is a well-known example of a galaxy undergoing ram pressure stripping (e.g. Kenney & Koopmann, 1999; Kenney et al., 2004; Vollmer et al., 2004; Chung et al., 2009). Furthermore, the direction of the stripping appears to be aligned along the direction towards BC6. However, BC6 and NGC 4522 are separated by 1800 km/s along the line-of-sight. Vollmer et al. (2006) attempted to simulate the gas stripping occurring in NGC 4522, finding that it was best fit by a very high velocity relative to the ICM (3500 km s), with a velocity component in the SE direction of 3000 km s. At this velocity the distance between NGC 4522 and BC6 could be traversed in 100 Myr. Vollmer et al. (2006) estimate that the period of maximum stripping pressure on NGC 4522 was approximately 50 Myr ago. These two time frames do not exactly line up, but Vollmer et al. (2006) indicated that significant stripping is expected to begin well before the peak ram pressure is reached. With these high velocities it may also be possible to explain the large velocity offset between BC6 and NGC 4522, if there is a significant component of NGC 4522’s velocity (relative to the ICM) along the line-of-sight.
Another point in favor of NGC 4522 is that ram pressure stripping simulations (e.g. Tonnesen & Bryan, 2012; Lee et al., 2022) show that most SF in stripped gas occurs either in the immediate vicinity of the galaxy or in the distant extremity of the tail. As BC6 is the only apparent stellar counterpart of the entire H i complex (see Appendix A) and is at the NW end of the tail, to be consistent with these simulations we would expect the point of origin to be to the SE, which NGC 4522 is.
Minchin et al. (2019) consider NGC 4522 as a possible progenitor, but reject it on a couple of grounds. Firstly, they argue that if the H i complex originated from NGC 4522 then you might expect cloud C5 to be at higher radial velocity than C3 (cf. Figure 3), because it is closer to NGC 4522, but the reverse is true. However, the velocity dispersion of the whole structure is quite low and such an inconsistency might be explained by the two clouds being stripped from different sides of the galaxy (thus having different initial velocities). Also, the initial velocity relative to the ICM would be rapidly reduced, thus the original objection may not be valid if the H i complex has been separated from NGC 4522 for long enough for the entire structure to come to rest relative to the ICM.
The second issue raised by Minchin et al. (2019) is that it is unlikely that the local ICM radial velocity in this region of Virgo is as low as 500 km s (recall that in §2 we mentioned that the typical velocity of galaxies in this regions of Virgo is 1000 km s). If this is not the case then it is not plausible that the H i gas could have initially had a radial velocity of 2300 km s and then slowed to 500 km s. This is a more serious objection to NGC 4522 as the progenitor and is difficult to dismiss. However, we note that a similar (though less severe) objection applies to all the candidate progenitor galaxies to the NW and SE listed in Table 3, as these all have radial velocities above 500 km s. If one of these is indeed the origin of BC6, then this may be an indication that we do not understand the motions (and perhaps, sub-structure) of the ICM in this region of Virgo.
Finally, Minchin et al. (2019) noted that the apparent gap in H i between the ram pressure tail still connected to NGC 4522 and the H i complex around BC6 would require two stripping episodes, with a break in between. However, simulations such as Kapferer et al. (2009) and Tonnesen & Bryan (2012) show similar morphology, with significant gaps in the H i column density of tails despite continuous stripping. Gas that is initially heated (and therefore not visible in 21 cm line emission) may cool and condense back into H i clouds in the tail, especially if the gas is metal-rich (e.g. Lee et al., 2022). However, it is not clear if this is a viable mechanism for such a large quantity of gas as is associated with BC6.
To summarize, we have considered a number of potential progenitor galaxies within a few degrees of BC6, but there is no clear parent object, in fact, all the candidates appear to have at least one disqualifying factor. We consider NGC 4522 to be the most likely parent object because it has similar metallicity and is a known case of extreme ram pressure stripping with the gas tail pointing towards BC6. However, it is difficult to account for the kinematic separation of NGC 4522 and BC6. NGC 4380 and NGC 4519 also remain candidates for the progenitor, but this may simply be because there are fewer observations of these galaxies and we could be ignorant of disqualifying information. NGC 4445 was suggested as the point of origin by Kent et al. (2009) and Minchin et al. (2019), but we disfavor this as Sorgho et al. (2017) seem to show that it has a gas tail that points away from BC6 and because the observed H i complex represents almost the entirety of its estimated original gas content.
5.2 Fate
Here we briefly consider the evaporation timescale for the clouds seen in the H i complex, as a means to determine the fate of BC6. Following the discussion of Borthakur et al. (2010), which itself is based on Cowie & McKee (1977) and Vollmer et al. (2001), we see that cloud evaporation can proceed in either the classical or saturation scenario depending on the value of , the ratio of the classical and saturated heat flux through the spherical shell of the cloud, where
and is the temperature of the hot ICM, is its number density (in cm), and is the cloud radius (in pc). To evaluate this and proceed we need to know the physical properties of both a cloud within the H i complex and the surrounding ICM.
To obtain approximate physical properties of a single cloud (as far as possible) we take the values for C2 (Figure 3, left) which was observed with the VLA in Kent et al. (2009). In this higher resolution imaging the cloud has an elongated structure with two main concentrations of mass. For simplicity we take a quarter of the H i major axis (1.5′) measured as the radius of a single cloud (2 kpc) and half of the total H i mass of C2 as its mass ( M). Here we have also multiplied by 1.4 to account for Helium and other metals. Assuming a spherical geometry gives a cloud number density of 0.3 .
For the ICM properties we use the simple -model ICM density profile used in Vollmer et al. (2001), evaluated at 2.7 (the separation between BC6 and M 87), to obtain a ICM density estimate of cm. For the temperature we see from the profile plotted by Böhringer et al. (1994) that at 2.7 from M 87 the average temperature of the ICM is roughly 2 keV ( K).
Using the above values in Equation 1, we see that this cloud is in the saturated regime (i.e. ) where the evaporation speed is limited by the rate heat can be conducted through the shell of the cloud. The expression derived by Borthakur et al. (2010) for the lifetime of an H i cloud in this regime is
where is the average number density of the cloud (in cm). The other quantity are the same as in Equation 1. Evaluating this expression gives a cloud lifetime for C2 of over a Gyr.
Although this is only an order of magnitude estimate it is clear from the stellar population (§4.1) that the H i complex has been traversing the ICM for at least 100 Myr and is still an extremely rich H i structure. Furthermore, Calura et al. (2020) considered in detail the survival of a much smaller gas cloud M (designed to mimic SECCO 1), finding that when simulating it traversing the ICM at high speed it could still survive as a distinct structure for around a Gyr. Thus, given the enormous reservoir still present, it seems that BC6 is not in imminent danger of losing its H i gas and is unlikely to become unbound and disperse into the cluster in the short term.
However, as the H i complex will not have a significant dark matter content, it unquestionably cannot survive indefinitely in the Virgo cluster. On long timescales it will eventually evaporate and its aging stars will disperse and contribute to the intra-cluster light. Indeed, a Gyr is on the order of the orbital timescale of the cluster, thus it is possible that the complex would undergo interactions with inhomogeneities in the ICM within that time or, given its scale, might experience a strong tidal interaction with a galaxy. Either of which could act to disperse it and shorten its expected lifetime. In addition, if SF continues then stellar feedback may also act to disperse the structure, especially as it is not supported by the additional potential well of a dark matter halo. Thus, the evaporation timescale is likely an upper limit on how long this structure can survive, even if individual clouds might in theory survive that long.
6 Conclusions
We have identified the stellar counterpart (BC6) of an enormous, isolated H i structure (190 kpc long, M) in the Virgo cluster that has been thought to be “dark” since its discovery by Kent et al. (2007). The extremely low (stellar) mass counterpart lies at the NW tip of the structure (Figure 3). In ground-based imaging it appears as faint and irregular blue clumps, which HST images reveal to consist exclusively of young (100 Myr) stars. Given this population, the total stellar mass of BC6 is likely only a few times M, making it a truly extreme object in terms of gas fraction.
An optical spectrum of this counterpart from HET returned an H redshift (500 km s) consistent with the velocity of the co-spatial H i gas seen in ALFALFA and with GBT observations, thereby confirming association. This spectrum also revealed that BC6 is metal-rich (given its low stellar mass), indicating that its gas complex must have originated in a galaxy where the gas could be pre-enriched.
Given these properties, BC6 appears to be a member of the novel stellar systems described by Jones et al. (2022b), known as “blue blobs,” standing out only because of its extraordinary H i content. The origin of this gas is as challenging to conclusively determine as are the origins of previously studied blue blobs. However, high speed ram pressure stripping remains the most plausible explanation, regardless of which cluster member was the origin. We favor NGC 4522 as a candidate point of origin as it has a similar metallicity, is a known example of ongoing extreme ram pressure stripping, and probably passed through the local peak of X-ray emission (from the ICM), where BC6 currently resides, about 100 Myr ago (Böhringer et al., 1994; Vollmer et al., 2001). However, this association is far from conclusive, in particular the kinematic separation of BC6 and NGC 4522 is difficult to explain.
Given how long this H i structure was known without a stellar counterpart being discovered, it seems likely other “dark” H i structures in Virgo (e.g. Kent et al., 2007; Taylor et al., 2012) may host similarly faint blue blob counterparts as well. This is a possibility that we aim to investigate through our ongoing citizen science search covering the entire cluster.
Appendix A Other counterpart candidates
Although our citizen science search for additional blue blob candidates throughout Virgo is incomplete at the time of writing, we extracted all the candidates that currently exist in this region to identify if BC6 is the only object associated with the H i tail, or if there might be other candidate blue blobs coincident with other clouds.
We only found one additional candidate (Figure 6) coincident with the H i structure, next to cloud C4 (Figure 3). The appearance of this candidate suggests that it is likely a background group of interacting galaxies. Although there are blue clumps, as in genuine blue blobs, they are surrounded by diffuse yellow light, probably made up of old stars that are very far away. Thus, we do not consider this likely to be a genuine blob blue or associated with the H i structure, however, we plan to obtain optical spectroscopic follow-up in the future to rule this out.